OpenStudy (anonymous):

Log2 x + 5 = 8 - log2 (x + 7)

OpenStudy (anonymous):

add log2 (x + 7) to both sides, substract 5 from both sides, rewrite the log expression log2 x + log2 (x+7)as the log of a product. exponentiate both sides. get x

OpenStudy (anonymous):

u should get x = -1 and x = 8, but since x = -1 is outside the domain of the logarithm, reject it

OpenStudy (anonymous):

can u show me the workk but i think im doing it completely wrong...heres what i came up with...lol log2 x + log2 (x+7) = 3 log2 [x/(x+7)] 3 [x/(x+7)] = 6 x= 6(x+7)

OpenStudy (anonymous):

ok for step 2, log2 (x(x+7)) = 3, exponentiate both sides using 2, you will get x(x+7) = 8, then by collecting terms, x^2 - 7x - 8 = 0, solve for x

OpenStudy (anonymous):

u should get x = 8 and x = -1, but since the log is not defined at x = -1, you must reject this answer

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

ohhhh okayy so u get x=8 right??

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

:)